Lovely123221:

A boy has 11 coins in dimes and quarters. Their value is $1.70. How many of each does he have? Which of the following equations could be used to solve the problem? x (11 - x) = 1.70 25x 10(11 - x) = 1.70 0.10x 0.25(11 - x) = 1.70

Babyplier:

@mercury

Mercury:

I think there are supposed to be + signs where the big spaces are? x + (11 - x) = 1.70 25x + 10(11 - x) = 1.70 0.10x + 0.25(11 - x) = 1.70 anyway, we want to write an equation that expresses the total value of coins in terms of x if we let x = # of dimes, notice how 0.10*x gives the monetary value of x dimes write a similar expression for the # of quarters, then add them together and sum to the total value. the result should match one of the three equations.

Lovely123221:

Thank you so much! And yes the spaces were supposed to be plus signs. :)